The equations of 2 lines are shown below:
2x-4y=6
3x+y=-5
What are the coordinates of the point of intersection?

Found 2 solutions by checkley71, venugopalramana:
Answer by checkley71(8403)

2X-4Y=6 3X+Y=-5 MULTIPLYING EQUATION 2 BY 4 WE GET 12X+4Y=20 ADDING TO EQUATION
12X+4Y=-20
2X-4Y=6
-------------
14X=-14 OR X=-1 & 2(-1)-4Y=6 OR -2-4Y=6 OR -4Y=8 OR Y=8/-4 OR Y=-2

Answer by venugopalramana(3286)

lines are shown below:
2x-4y=6...................................I
3x+y=-5.............................II
What are the coordinates of the point of intersection?
WE HAVE TO SOLVE THE 2 EQNS.AS THE POINT OF INTERSECTION SHOULD LIE ON BOTH LINES AND HENCE SHOULD SATISFY BOTH EQNS.
EQN.I+EQN.II*4..GIVES
2X-4Y+12X+4Y=6-20=-14
14X=-14
X=-1................SUBSTITUTING IN EQN.II
3*(-1)+Y=-5
-3+Y=-5
Y=-5+3=-2
HENCE POINT OF INTERSECTION IS (-1,-2)